A monopolist faces two consumer groups: old and young. The inverse demand of old clients for the output of the monopolist is Po = 130 – 3Qo.  The inverse demand of young clients for the output of the monopolist is Py = 300 – 6Qy.  The marginal cost of supplying any type of client is MC = 10.  If the monopolist can price discriminate between the two groups

  1. What will be the quantity for old?

Select one:

a.

  1. 0o <10;
  1. 0o <10;

0o <10;

  1. 0o <10;
  1. 0o <10;
  1. 0o <10;

b.

  1. 10o <30;

c.

  1. 30o <60;

d.

  1. 60o <100;

what price will old and clients be charged?

Select one:

a.

  1. 0<Po <10;

b.

  1. 10<Qo <30;

c.

  1. 30<Qo <60;

d.

  1. 60<Qo <100;

e.

  1. None of the above

What will be the quantity for young?

Select one:

a. 0<Qy <10

b.

  1. 10<Qy <30;

c.

  1. 30<Qy <60;

d.

  1. 60<Qy <100;

e. None of the above

what price will young clients be charged?

Select one:

a.

  1. 0<Py <10;

b. 10<Py <30

c.

  1. 30<Py <60;

d.

  1. 60<Py <100;

e.

  1. None of the above

What will be the firm’s profit?

Select one:

a.

  1. 0<pi <1000;

b.

  1. 1000<pi<3000;

c.

  1. 3000<pi <6000;

d.

  1. 6000<pi <10000;

e.

  1. None of the above
Answer #1

Prost group: P = 30-300 Mc =10 MR TRo 1300-802 MRO = 130 - 600. In qulbrun , MRO - MG 130 - 690 = 10 120 = 600 20 = Co Po=130let me answer here 1. (c) 3Qo<60 (old quantity) 2. (d) Po 8 Go LP <loo (old price) 3 (6) 10LQy (30 4. Py = 156 (e) none of th